How do i find the ordered pairs for x-6=0 with y =1, 4, -5
Answer by Edwin McCravy(20054)

How do i find the ordered pairs for x-6=0 with y =1, 4, -5

x-6 = 0  just means x = 6,  so you just put the x-coordinate 6
with each of those values of y in an ordered pair, so you have
these three ordered pairs:

(6,1) by putting the x coordinate 6 with the y-coordinate 1
(6,4) by putting the x coordinate 6 with the y-coordinate 4
(6,-5) by putting the x coordinate 6 with the y-coordinate -5

You see, there is nothing to calculate in this special case,
you just make the ordered pairs.

Then when you plot those three points, you have this:



Then you get a ruler and draw a straight line through them, like the
green one below:



That's the graph of the equation x-6 = 0

Edwin